You separated your Php500,000 into two accounts. The First account is placed at a rate of 111% while the second at a rate of 6%. The annual interest of the first account is Php1,210 more than the annual interest of the second account. How much is invested at 6%?
Answer by checkley77(12844)

1.11x=.06(500,000-x)+1,210
1.11x=30,000-.06x+1,210
1.11+.06x=31,210
1.17x=31,210
x=31,210/1.17
x=$26,675.21 amount invested @ 111%.
500,000-26,675.21=$473,324.79 amount invested @ 6%.
Proof:
1.11*26,675.21=.06*473,324,79+1,210
29,609.48=28,399.48+1,210
29,609.48=29,609.48